# Collimated $\gamma$-flash emission along the target surface irradiated by a laser at non-grazing incidence

### v1

@@ 4.1 Generation of collimated $\gamma$-photon beam along the target surface
The ${}^{\circ}$incidence case produces approximately one order of magnitude more $\gamma$-photons compared to the normal incidence case. The $\gamma$-photon cut-off energy for the ${}^{\circ}$incidence case is above 350 MeV, compared to less than 300 MeV for the normal incident case (the bin size is 5 MeV and the minimal value of the dN/dE = 284160 corresponds to five macroparticles in the ${}^{\circ}$case). The inset of Fig. LABEL:fig:spectra_rate-a shows the angular distribution of $\gamma$-photon energy in the $p_{y}$/$p_{x}$ plane, using $\gamma$-photons fulfilling eq. 6 as discussed above. The ${}^{\circ}$incidence simulation predominantly emits $\gamma$-photons in the direction parallel to the target surface (at $90^{\circ}$, depicted by the black line), while the normal incidence case is characterized by the typical double-lobe pattern. Moreover, the ${}^{\circ}$incidence case emits approximately 3 times more $\gamma$-photons at the ${}^{\circ}$emission angle compared to the maximum of the normal incidence case. The time evolution of electron energy spectra is shown in fig. LABEL:fig:spectra_rate-b. The cut-off electron energy is almost double compared to the $\gamma$-photon energy, approaching 700 MeV. The number of high-energy electrons with energy above 300 MeV drops significantly between 65 fs and 70 fs when a high number of $\gamma$-photons is emitted. The angular distribution of electron energy is shown in the inset of Fig. LABEL:fig:spectra_rate-b. Here, only electrons with $\chi_{e}>0.1$ are considered in the distribution. At the time of 60 fs the high $\chi_{e}$ electrons have their momentum oriented at ${}^{\circ}$(target surface). At later times, the momentum orientation of these electrons gradually drift away as the target is bended by the radiation pressure. Therefore, a broadening of the $\gamma$-photon angular energy distribution gradually occurs, as seen in the inset of Fig. LABEL:fig:spectra_rate-a. At a time of 70 fs, the electron angular distribution separates into several directions. However, by this time their effect on the $\gamma$-photon distribution is becoming less important, as the total number of electrons with high $\chi_{e}$ is decreasing.
@@ 4.2 Dependence on the laser incidence angle
In this Section we compare simulations with incidence angles ranging from 30 to 60. As predicted in Ref. [70], the maximum energy of electrons accelerated in the interference pattern increases with the angle. This behavior is shown in Fig. LABEL:fig:angular_distribution_angles-a at time of 65 fs, i.e., when the electron acceleration in the interference pattern is still ongoing. At a later time, the electrons are directly accelerated by the reflected laser pulse in the specular direction (see Ref. [81] and the references cited therein). However, the directionality of the electron beams gradually shifts away from the target surface when larger angles are used, as shown in the inset of Fig. LABEL:fig:angular_distribution_angles-a, where we consider electrons of $\chi_{e}>0.1$.
The $\gamma$-photon energy spectra at 100 fs are shown in Fig. LABEL:fig:angular_distribution_angles-b. Both the number of emitted $\gamma$-photons and their cut-off energy increase with the angle, although this increase is less apparent than that of the electrons. Exceptionally, for the ${}^{\circ}$incidence angle case more $\gamma$-photons are emitted than in the ${}^{\circ}$case. As seen in the inset of Fig. LABEL:fig:angular_distribution_angles-a the ${}^{\circ}$case develops an additional central lobe at ${}^{\circ}$and the spectrum contains higher energy density of electrons with $\chi_{e}>0.1$ than the ${}^{\circ}$case. Consequently, the $\gamma$-photon energy angular distribution (inset of Fig. LABEL:fig:angular_distribution_angles-b) of the ${}^{\circ}$case contains more energy than the ${}^{\circ}$case. Therefore, the interference pattern effect, described in Section 4.1, is utilized more efficiently in the ${}^{\circ}$incident angle case than in the ${}^{\circ}$case.
Overall, the angular dependency of the $\gamma$-photon density distribution (inset of Fig. LABEL:fig:angular_distribution_angles-b) shows a gradual shifting from the double-lobe structure, for the normal incidence case, to a single lobe along the target surface for the ${}^{\circ}$case. The ${}^{\circ}$and ${}^{\circ}$cases reveal a smooth transition between the ${}^{\circ}$and ${}^{\circ}$cases. When increasing the incident angle above 45, a third lobe is emitted near the laser specular direction. In addition, all our simulations exhibit a local minimum approximately in the direction of the incident laser pulse.
@@
The laser energy (283 J) to $\gamma$-photon energy conversion efficiency for various incidence angles is shown in Table 1. The simulations with oblique incidence converted up to 3.5 times more energy into $\gamma$-photons than the simulation with normal incidence. The conversion efficiency rises with the angle till 52.5, where it reaches 10.7 %, and decreases afterwards. The rest of Table 1 shows the maximum $\gamma$-photon energy reached in the simulations (second line) and cut-off $\gamma$-photon energies at dN/dE = 284160 (third line), shown in Fig. LABEL:fig:angular_distribution_angles-b.
Table 1: The conversion efficiency of the laser energy into $\gamma$-photon energy, the maximum reached $\gamma$-photon energy and the cut-off $\gamma$-photon energy at dN/dE = 284160 (the lower limit in Fig. LABEL:fig:angular_distribution_angles).
@@
4.3 Dependence on laser polarization
In this section the reference case of 45 with the linearly p-polarized laser is compared to simulations with the linearly s-polarized and circularly polarized (c-polarized) lasers. Note that c-polarization case behaves as an elliptical one prolonged in one direction on the target surface, as oblique incidence is used.
Unlike the compact structure in the energy distribution for the p-polarization case (Fig. LABEL:fig:spherical_polarizations-a), the energy distribution for the s-polarization case is spread over ${}^{\circ}$outside of the incidence plane (Fig. LABEL:fig:spherical_polarizations-b). Therefore, the maximum energy density in the s-polarization case is about an order of magnitude lower than in the p-polarization case. In the plane of incidence, the spreading of the emitted $\gamma$-photons is significantly less than in the other plane. Fig. LABEL:fig:spherical_polarizations-d shows the angular energy distribution around the incidence plane (eq. 6). Before normalization, the data for the s-polarization and c-polarization were 12.6 and 38.7 times smaller than the p-polarization, respectively. From Fig. LABEL:fig:spherical_polarizations-d one can deduce that the $\gamma$-photon spreading in the $x$-$y$ plane is minimal (as opposed to the other direction) for the s-polarization case and maximal for the c-polarization case.
The energy distribution of the c-polarization case (Fig. LABEL:fig:spherical_polarizations-c) spreads significantly in both directions. In the incidence plane three distinct spots are visible approximately at ${}^{\circ}$(incidence angle), ${}^{\circ}$(surface direction) and ${}^{\circ}$(specular angle). As in the case of s-polarization, the energy density distribution spreading results in a much lower peak value (38.7 times) compared to the p-polarization case.
4.4 Dependence on laser pulse duration, laser power and target thickness
In this section the reference case of 45 is compared to simulations with the same incident angle but with different laser pulse duration, power and target thickness.
Firstly, we compare simulations of different laser pulse durations (Fig. LABEL:fig:power_x_thickness-a), but keeping the same intensity. As expected, longer pulses (larger energy) emit more $\gamma$-photons. The dominant central lobe along the target surface occurs for all tested pulse durations. For a 15 fs laser pulse, the angular spread is lower than the reference case of 20 fs. On the contrary, for longer pulses, an additional lobe develops close to the specular angle. This behavior resembles that of increasing the incident angle over ${}^{\circ}$(inset of Fig. LABEL:fig:angular_distribution_angles-b) and is ascribed to target bending (by the rising part of the laser pulse), which changes the actual incidence angle. The target bending is mitigated for lasers with lower intensity.
Secondly, the effect of the target thickness is presented in Fig. LABEL:fig:power_x_thickness-b. The figure exhibits minimal changes for target thicknesses ranging between 0.5 and 2.5 $\mathrm{\mu m}$. For those thicknesses the iron target is sturdy enough to prevent the 14 PW laser from burning through the target, before the main $\gamma$-photon emission occurs. Decreasing the target thickness to 0.252 $\mathrm{\mu m}$ results in the pulse to penetrate the target, thus limiting the $\gamma$-photon emission.
Lastly in Fig. LABEL:fig:power_x_thickness-c, we compare simulations with different laser power (intensity), while keeping the pulse duration and focal spot the same. The laser power ranges from 1.4 PW to 70 PW (peak intensity from $2\times 10^{22}\ \mathrm{W/cm^{2}}$ to $10^{24}\ \mathrm{W/cm^{2}}$). The values of the figure are normalized. The beam profile of 7 PW case is similar to the 14 PW one, although the maximum energy density is 4.7 times lower. In the 1.4 PW case the beam is even more collimated. However, the maximum energy density is reduced by 300 times. The 70 PW case provides the most collimated beam and the maximum energy density is 23.5 times larger compared to the 14 PW case. At the end of the simulation the laser pulse penetrates the target, as is shown in Fig. LABEL:fig:power_x_thickness-d. Therefore, the $\gamma$-photon yield can be increased with the use of thicker targets.
@@ 5 Conclusion
The best $\gamma$-photon beam collimation is obtained for ${}^{\circ}$incident angle with brilliance approaching $10^{23}$ s-1mm-2mrad-2 per 0.1% bandwidth, which is similar to that of Ref. [62]. The collimated patterns prevails in the incidence angle range of ${}^{\circ}$and 52.5. The laser energy to $\gamma$-photon energy conversion efficiency reaches up to 10.7 %. Although the use of a larger incident angle results in a higher electron energy, the laser energy conversion efficiency into $\gamma$-photons and $\gamma$-photon beam collimation significantly decrease as the angle approaches 60.
The collimated $\gamma$-photon beam occurs when a p-polarized laser is used. The use of s-polarization results in spreading of the $\gamma$-photon emission distribution in the laser polarization direction and the maximal energy density is lower by an order of magnitude. The use of circular polarization provides an even broader distribution, thus significantly reducing the $\gamma$-photon beam brightness.
A similar angular spread of the $\gamma$-photon beam is observed for laser powers ranging from 1.4 PW to 70 PW, for our simulation parameters. For target thicker than 0.5 $\mathrm{\mu}$m the $\gamma$-photon beam angular profile is almost identical. For thinner targets the 14 PW laser penetrates the target, resulting in decrease of the $\gamma$-photon emission.
Increasing the laser pulse duration or the 14 PW case results in the formation of an additional $\gamma$-photon beam in the specular direction. Nevertheless, the $\gamma$-photon beam emitted parallel to the target surface is dominant for the laser pulse durations up to 30 fs in our simulations. The effect is ascribed to the target bending by the lower intensity part of the laser pulse before the interaction with the peak of the laser pulse happens. Thus, significantly increasing the laser pulse duration results in a similar effect to the use of larger incidence angles for lasers with comparable power and other parameters.

### v2

@@ 4.1 Generation of collimated $\gamma$-photon beam along the target surface
The ${}^{\circ}$incidence case produces approximately one order of magnitude more $\gamma$-photons compared to the normal incidence case. The $\gamma$-photon cut-off energy for the ${}^{\circ}$incidence case is above 350 MeV, compared to less than 300 MeV for the normal incident case (the bin size is 5 MeV and the minimal value of the dN/dE = 284160 MeV-1 corresponds to five macroparticles in the ${}^{\circ}$case). Fig. LABEL:fig:spectra_rate-c shows the angular distribution of $\gamma$-photon energy in the $p_{y}$/$p_{x}$ plane, using $\gamma$-photons fulfilling eq. 6 as discussed above. The ${}^{\circ}$incidence simulation predominantly emits $\gamma$-photons in the direction parallel to the target surface (at $90^{\circ}$, depicted by the black line), while the normal incidence case is characterized by the typical double-lobe pattern. Moreover, the ${}^{\circ}$incidence case emits approximately 3 times more $\gamma$-photons at the ${}^{\circ}$emission angle compared to the maximum of the normal incidence case. The time evolution of electron energy spectra is shown in fig. LABEL:fig:spectra_rate-b. The cut-off electron energy is almost double compared to the $\gamma$-photon energy, approaching 700 MeV. The number of high-energy electrons with energy above 300 MeV drops significantly between 65 fs and 70 fs when a high number of $\gamma$-photons is emitted. The angular distribution of electron energy is shown in Fig. LABEL:fig:spectra_rate-d. Here, only electrons with $\chi_{e}>0.1$ are considered in the distribution. At the time of 60 fs the high $\chi_{e}$ electrons have their momentum oriented at ${}^{\circ}$(target surface). At later times, the momentum orientation of these electrons gradually drift away as the target is bended by the radiation pressure. Therefore, a broadening of the $\gamma$-photon angular energy distribution gradually occurs, as seen in Fig. LABEL:fig:spectra_rate-d. At a time of 70 fs, the electron angular distribution separates into several directions. However, by this time their effect on the $\gamma$-photon distribution is becoming less important, as the total number of electrons with high $\chi_{e}$ is decreasing.
@@ 4.2 Dependence on the laser incidence angle
In this Section we compare simulations with incidence angles ranging from 30 to 60. As predicted in Ref. [73], the maximum energy of electrons accelerated in the interference pattern increases with the angle, as can be seen in Figs LABEL:fig:angular_distribution_angles-a,b. The solid lines in Fig. LABEL:fig:angular_distribution_angles-a show the maximum electron energy at different simulation times. The dashed lines show the electron energy predicted by Eq. 5 (divided by a factor of two), which is in good agreement with our simulations. The model correctly predicts the increase of the maximum electron energy with incident angle. Fig. LABEL:fig:angular_distribution_angles-b shows the energy spectra at time of 65 fs, i.e., when the electron acceleration in the interference pattern is still ongoing. At a later time, the electrons are directly accelerated by the reflected laser pulse in the specular direction (see Ref. [84] and the references cited therein). However, the directionality of the electron beams gradually shifts away from the target surface when larger angles are used, as shown in Fig. LABEL:fig:angular_distribution_angles-c, where we consider electrons of $\chi_{e}>0.1$.
The $\gamma$-photon energy spectra at 100 fs are shown in Fig. LABEL:fig:angular_distribution_angles-d. Both the number of emitted $\gamma$-photons and their cut-off energy increase with the angle, although this increase is less apparent than that of the electrons. Exceptionally, for the ${}^{\circ}$incidence angle case more $\gamma$-photons are emitted than in the ${}^{\circ}$case. As seen in Fig. LABEL:fig:angular_distribution_angles-c the ${}^{\circ}$case develops an additional central lobe at ${}^{\circ}$and the spectrum contains higher energy density of electrons with $\chi_{e}>0.1$ than the ${}^{\circ}$case. Consequently, the $\gamma$-photon energy angular distribution (Fig. LABEL:fig:angular_distribution_angles-e) of the ${}^{\circ}$case contains more energy than the ${}^{\circ}$case. Therefore, the interference pattern effect, described in Section 4.1, is utilized more efficiently in the ${}^{\circ}$incident angle case than in the ${}^{\circ}$case.
Overall, the angular dependency of the $\gamma$-photon density distribution (Fig. LABEL:fig:angular_distribution_angles-e) shows a gradual shifting from the double-lobe structure, for the normal incidence case, to a single lobe along the target surface for the ${}^{\circ}$case. The ${}^{\circ}$and ${}^{\circ}$cases reveal a smooth transition between the ${}^{\circ}$and ${}^{\circ}$cases. When increasing the incident angle above 45, a third lobe is emitted near the laser specular direction. In addition, all our simulations exhibit a local minimum approximately in the direction of the incident laser pulse.
@@
The laser energy (283 J) to $\gamma$-photon energy conversion efficiency for various incidence angles is shown in Table 1. The simulations with oblique incidence converted up to 3.5 times more energy into $\gamma$-photons than the simulation with normal incidence. The conversion efficiency rises with the angle till 52.5, where it reaches 10.7 %, and decreases afterwards. The rest of Table 1 shows the maximum $\gamma$-photon energy reached in the simulations (second line) and cut-off $\gamma$-photon energies at dN/dE = 284160 MeV-1 (third line), which is the lower limit in Fig. LABEL:fig:angular_distribution_angles-d.
Table 1: The conversion efficiency of the laser energy into $\gamma$-photon energy, the maximum reached $\gamma$-photon energy and the cut-off $\gamma$-photon energy at dN/dE = 284160 MeV-1 (the lower limit in Fig. LABEL:fig:angular_distribution_angles-d).
@@ 5 Conclusion
The collimated patterns prevails in the incidence angle range of ${}^{\circ}$and 52.5. The laser energy to $\gamma$-photon energy conversion efficiency reaches up to 10.7 %. Although the use of a larger incident angle results in a higher electron energy, the laser energy conversion efficiency into $\gamma$-photons and $\gamma$-photon beam collimation significantly decrease as the angle approaches 60. The best $\gamma$-photon beam collimation is obtained for ${}^{\circ}$incident angle with brilliance approaching $10^{23}$ s-1mm-2mrad-2 per 0.1% bandwidth, which is similar to that of Ref. [63].
@@
Appendix A: Dependence on laser polarization, pulse duration, target thickness, preplasma, laser power, and Virtual Reality visualization
Firstly, we show the advantage of using p-polarization (p-pol) over s-polarization (s-pol) and circular polarization (c-pol) in our scheme. Note that for oblique incidence the circularly polarized laser does not interact with the plasma so symmetrically as in the normal incidence case. Fig. LABEL:fig:spherical_polarizations-a shows the angular energy distribution around the incidence plane (eq. 6) and reveals that the maximum values for p-pol are more than an order of magnitude larger than for s-pol (12.6 times) and c-pol (42.5 times). The beam profiles are normalized in the inset of Fig. LABEL:fig:spherical_polarizations-a. For further investigation we show the 3D angular energy distribution in Figs. LABEL:fig:spherical_polarizations-b,d,e). For the s-pol (Fig. LABEL:fig:spherical_polarizations-b) the energy distribution spreads over ${}^{\circ}$in the $\varphi$ direction, as it is the laser polarization plane. In the $\theta$ direction the distribution is around the incident angle of 45°. However, due to the significant spread in the $\varphi$ direction the maximum energy density is about an order of magnitude lower than in the p-pol case (see Fig. LABEL:fig:spherical_distribution_00x45-b). For the c-pol, the electric field interference structure in the x-y plane (similar to the p-pol shown in Fig. LABEL:fig:field_electron_gamma(b) but about 1/3 weaker) is shown in Fig. LABEL:fig:spherical_polarizations-c. Therefore, part of the $\gamma$-photons are generated along the target surface ($\theta=90$°) in a way similar to that described in Section 4.1 for the p-pol, although more spread in both directions as seen in Fig. LABEL:fig:spherical_polarizations-d. As the mechanism along the target surface is weaker for c-pol, the secondary beam (between 0° and 45°) reaches similar values as the main one (around 90°). Fig. LABEL:fig:spherical_polarizations-e) shows the $\gamma$-photon energy distribution generated between 75 fs and 100 fs, when the laser pulse is reflected. During this interaction a third beam is generated around the specular angle of 135°.
Secondly, we compare simulations of different laser pulse durations (Fig. LABEL:fig:power_x_thickness-a), but keeping the same intensity. As expected, longer pulses (larger energy) emit more $\gamma$-photons. The dominant central lobe along the target surface occurs for all tested pulse durations. For a 15 fs laser pulse, the angular spread is lower than the reference case of 20 fs. On the contrary, for longer pulses, an additional lobe develops close to the specular angle. This behavior resembles that of increasing the incident angle over ${}^{\circ}$(LABEL:fig:angular_distribution_angles-e) and is ascribed to target bending (by the rising part of the laser pulse), which changes the actual incidence angle. The target bending is mitigated for lasers with lower intensity.
Thirdly, the effect of the target thickness is presented in Fig. LABEL:fig:power_x_thickness-b. The figure exhibits minimal changes for target thicknesses ranging between 0.5 and 2.5 $\mathrm{\mu m}$. For those thicknesses the iron target is sturdy enough to prevent the 14 PW laser from burning through the target, before the main $\gamma$-photon emission occurs. Decreasing the target thickness to 0.252 $\mathrm{\mu m}$ results in the pulse to penetrate the target, thus limiting the $\gamma$-photon emission.
Fourthly, the effect of the preplasma is presented in Fig. LABEL:fig:power_x_thickness-c, since it was shown that an appropriate preplasma can reduce the divergence angle of the electron beam [85] and thus influence the $\gamma$-photon generation. We reduced the target thickness to 1 $\mathrm{\mu m}$ and introduced a 1.75 $\mathrm{\mu m}$ preplasma in front of it (positions $x<0$) with density ranging from 0.0013 $n_{c}$ to the target density. The preplasma profile is $n_{e}\exp(-|x|/L)$ with the preplasma scale length $L=0.127$ $\mathrm{\mu m}$. The value of $L$ is chosen to follow the condition $L=c/\omega$. This condition is beneficial for generation of high-order harmonics [86]. The angular spectra shows a rise in the direction along the target surface by about 23%.
Lastly in Fig. LABEL:fig:power_x_thickness-d, we compare simulations with different laser power (intensity), while keeping the pulse duration and focal spot the same. The laser power ranges from 1.4 PW to 70 PW (peak intensity from $2\times 10^{22}\ \mathrm{W/cm^{2}}$ to $10^{24}\ \mathrm{W/cm^{2}}$). The values of the figure are normalized. The beam profile of 7 PW case is similar to the 14 PW one, although the maximum energy density is 4.7 times lower. In the 1.4 PW case the beam is even more collimated. However, the maximum energy density is reduced by 300 times. The 70 PW case provides the most collimated beam and the maximum energy density is 23.3 times larger compared to the 14 PW case. At the end of the simulation the laser pulse penetrates the target, as is shown in Fig. LABEL:fig:power_x_thickness-e. Therefore, the $\gamma$-photon yield can be increased with the use of thicker targets.
Additionally, the visualization of the $\gamma$-photon emission temporal evolution of the ${}^{\circ}$case can be accessed online [77] via our Virtual Beamline – VBL application [78, 79, 80]. VBL uses a custom-made WebGL [87] app to render the visualization inside a regular web browser, as shown in Fig. LABEL:fig:power_x_thickness-f. The application also enables support for Virtual Reality (VR) viewing, which was tested with Oculus Rift S and HP Reverb G2 headsets.